OpenStudy (anonymous):

Could use some help if someone is kind enough Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -1 x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

OpenStudy (anonymous):

use y = 1-2x and put it into the second equation 4x + 2(1-2x) = -1 solve this for x

OpenStudy (anonymous):

4x + 2(1-2x) = -1 8x+2=-1 -2 -2 8x=-3 x=-3/8 x=-0.375?

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (jack1):

heya man, these are parallel lines

OpenStudy (jack1):

so quick rule: there would be an infinite number of solutions if they were the same line (as they always meet, in loads of places) there would be no solutions if they were parallel lines (as they never meet)

OpenStudy (jack1):

does this help at all, do u want to know how they're parallel, u all good now? lemme know?

OpenStudy (anonymous):

so then its no solution?

OpenStudy (jack1):

aye, no solution, they have the same gradient but different y intercepts, so they're parallel and never meet

OpenStudy (anonymous):

ah i see

OpenStudy (anonymous):

thenks

OpenStudy (jack1):

welcome man, slaters ;)

OpenStudy (anonymous):

Good that @Jack1 noted these are parallel lines, with no solution if njot identical. Substitution went wrong when it was concluded that 4x + 2(1-2x) = 8x + 2, rather than 0x +2, which cannot =-1 and tells nothing about x.